40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 851     Entry time: Tue Feb 28 22:50:11 2012
Author: Frank 
Type: Notes 
Category: Schedule 
Subject: update 

The following tasks will be finished within the next 24h:

  • aligning both cavities
  • align beat
  • install RF-AM PD (start with one to get estimate for noise budget)
  • measure TF from seismic to beat (if possible, vacuum chamber temp might not be stable enough)

we will then focus on (re-)measuring some things we need very soon for a more accurate NB (there are more but the following ones are my favorites for now. We are about a factor of 3 above the CTN @200Hz)

  • RF-AM - not in NB yet, but seems to dominate 100Hz-1kHz region, so we do this first (seen last week while optimizing everything)
  • RIN  - old measurement must be wrong as it is too high
  • seismic  - new stack, so new TF
  • phase noise of Marconi locked to rubidium clock - we know that the noise will be less below 1kHz, above it most likely does not change the phase noise level
  • electronic noise from both PDH loops

 

ELOG V3.1.3-